OpenStudy (anonymous):

The perimeter of a rectangle is 256 inches. How long are the sides if the length exceeds the width by 16 inches? Mark lets x represent the width of the rectangle and writes the equation 2x + 2(x + 16) = 256. Solving for x, the width of the rectangle is A.56 B.60 C.68

OpenStudy (anonymous):

I ignored the equation. 56+16=72 72+72+56+56=256

OpenStudy (anonymous):

I'm not good with equations, so I just worked with the numbers until I found the answer. Maybe DemolisionWolf will show you the right way to do it.

OpenStudy (anonymous):

solve this: 2x + 2(x + 16) = 256 so, first step is use the Distributive Law 2x + 2x + 32 = 256 next, combine like terms 4x + 32 = 256 next, isolate x 4x + 32 -32 = 256 -32 4x = 224 4x/4 = 224/4 x= 224/4

OpenStudy (anonymous):

X=56

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Length: 72 Width: 56

OpenStudy (anonymous):

^_^